OpenStudy (anonymous):

Find the exact area of the circle with the given radius. r = 3 mm A. 9 mm2 B. 9π mm2 C. 3π mm2 D. 9.42 mm2

OpenStudy (anonymous):

the formula to find the area of a circle is \[A = \pi r^2 \]

OpenStudy (anonymous):

so b

OpenStudy (anonymous):

\[\pi 3^2 = ??\]

OpenStudy (anonymous):

yes, B

OpenStudy (anonymous):

thx